OpenStudy (anonymous):

What is the length of the minor axis of the ellipse given by the equation x squared over 144 plus y squared over 25 = 1?

OpenStudy (jdoe0001):

\(\bf \cfrac{x^2}{144}+\cfrac{y^2}{{\color{brown}{ 25}}}=1\implies \cfrac{x^2}{12^2}+\cfrac{y^2}{{\color{brown}{ 5}}^2}=1 \\ \quad \\ \quad \\ \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{{\color{brown}{ b}}^2}=1\qquad minor\ axis\to 2{\color{brown}{ b}}\)

OpenStudy (jdoe0001):



OpenStudy (anonymous):

thats not one of the answers for my homework

OpenStudy (anonymous):

wait is the answer 5 ?